A data set lats weights bj of plassic dascerded by households. The highest weight is 5.26 b, the mean of alil of the weights is 2.119 b, and the standard deviation ofthe weigt, s s-1058 t. a What is the diference between the weight of 5.26 lo and the mean of the weights? b. How many standard deviations is that the deerence found in part (alT c. Convert the weight of 5 26 ib to a z score d. f we consider data speeds that convent to a scores between -2 and 2 to be neither signriticanlty low nor signficanity high, is the weight or 5.26 b signvtcant? a. The diflerence is Type an integer or a decimal. Do not round ) ern grwer in the answer box and then click Check Answer
Answer #1

Answer:

a). difference: 5.26-2.119 = 3.141 lb

b). 3.141/1.058 = 2.9688 standard deviations

c).

z = (x-µ)/σ

z = (5.26-2.119)/1.058

=2.9688

Or

2.97( two decimals)

d).

z value 2.97 is >2,

The weight of 5.26 is significantly high.